# Detecting PowerShell-based Fileless Cryptojacking Attacks Using Machine Learning

### v1

@@ Abstract
[…]
[uc-it] organization=School of Information Technology, organization=University of Cincinnati, city=Cincinnati, state=OH, country=USA
[…]
[uc-ece] organization=Department of Electrical and Computer Engineering, organization=University of Cincinnati, city=Cincinnati, state=OH, country=USA
@@ 10 Submission Declaration
    The work described has not been published previously except in the form of a preprint, an abstract, a published lecture, academic thesis or registered report.
@@
  • 3.
    If accepted, the article will not be published elsewhere in the same form, in English or in any other language, including electronically, without the written consent of the copyright-holder.
@@ 11 Authorship
12 Declaration of Interests
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.

### v3

# Detecting Fileless Cryptojacking in PowerShell Using AST-Enhanced CodeBERT Models

@@ 1 Introduction
An example of the encoded PowerShell-based fileless cryptojacking script code screenshot.
@@
A Sample PowerShell-based fileless cryptojacking attacks workflow. The flow consists of the 10 stages.
@@ 3.3 Long Short-term Memory (LSTM)
The LSTM block architecture diagrams including the three gates and the activation functions.
@@ 5.5 Model Definition - CodeBERT
The diagram blocks of BiLSTM and the information flow between blocks.
@@ 10 Submission Declaration
    The work described has not been published previously except in the form of a preprint, and academic thesis.
@@ 11 Authorship
12 CRediT Roles
Said Varlioglu: Conceptualization, Conceptualization, Investigation, Formal analysis, Visualization, Validation, Writing – original draft. Nelly Elsayed: Conceptualization, Methodology, Supervision, Writing – original draft, Writing – review & editing. Murat Ozer: Validation, Data curation, Writing – review & editing. Zag ElSayed: Validation, Formal analysis, Writing – review & editing. John M. Emmert: Supervision, Project administration.
@@ 13 Funding
14 Disclosure statement
No potential competing interest was reported by the author(s).
@@ 16 Data availability statementt
17 Biographical note
Said Varlioglu (Ph.D.), Digital Forensics and Incidence Response Engineer. Recieved his Ph.D. at Information Technology from Univeristy of Cincinnati in 2025.
Nelly Elsayed (Ph.D.), Associate Professor, Leader and Founder of the Applied Machine Learning and Intelligence (AMLI) Lab at the University of Cincinnati.
Murat Ozer (Ph.D.), Associate Professor at the University of Cincinnati.
Zag ELSayed (Ph.D.), Assistant Professor at the University of Cincinnati and a Fellow Researcher at the Cincinnati CHildren’s Hospital Medical Center
John Emmert (Ph.D.), Professor in the Department of Electrical Engineering and Computer Science and Director of the NSF Center for Hardware and Embedded Systems Security and Trust (CHEST) IUCRC.